# Adjusting for Heavy Censoring and Double-Dipping to Compare Risk Stratification Abilities of Existing Models for Time to Diagnosis of Huntington Disease

### v1

@@ 2.2.2 Model 2: CAG-Age Product (CAP)
The CAP model was introduced by Zhang et al. (2011) Zhang2011 to predict a patient’s time to diagnosis given their CAG repeats and current age. It was developed using data from the PREDICT-HD study, restricted to patients with (i) a CAG repeats $\geq 36$ and (ii) DCL <4<4 at study entry. That is, patients needed to be at-risk for Huntington disease but not yet diagnosed at enrollment. After applying these criteria, there were $730$ patients. During the study, $137$ of these patients were clinically diagnosed, while the remaining $593$ ($81\%$) were not and thus had right-censored times of diagnosis.
@@
The CAP model is defined $\hbox{log}(X_{i})=\beta_{0}+\beta_{1}\text{Age}_{i}(\text{CAG}_{i}+\beta_{2})+\epsilon,$ where ${\boldsymbol{\beta}}=(\beta_{0},\beta_{1},\beta_{2})$ denote the intercept and effects of $\text{Age}_{i}\times\text{CAG}_{i}$ and $\text{Age}_{i}$. The error term $\epsilon$ is assumed to follow a logistic distribution with location =0=0 and scale $\sigma$. The linear predictor, $\beta_{0}+\beta_{1}\text{Age}_{i}(\text{CAG}_{i}+\beta_{2})$, can be interpreted as a log-risk score. The parameters ($\boldsymbol{\beta}$}, $\sigma$) are estimated via MLE, with a likelihood function that accounts for right censoring. Time to diagnosis $X_{i}$ is assumed to be conditionally independent of $C_{i}$ given CAG repeats and current age. The CAP model can be estimated using the survreg() function from the survival package in R with option dist = "loglogistic" Therneau2022.
@@ 2.2.3 Model 3: Multivariate Risk Score (MRS)
The MRS model was developed with similar methods and intended applications to CAP. Designed with observational studies in mind, the model predicts the time from enrollment to HD diagnosis Long2015. However, compared to Langbehn and CAP, MRS was developed using a much broader set of motor, imaging, cognitive, functional, psychiatric, and demographic covariates.
The MRS model was developed using data from PREDICT-HD, subsetting to patients with (i) CAG repeats $\geq 36$ and (ii) DCL <4<4 at baseline. With these criteria (same as those for the CAP model), the analytic dataset had a sample size of $1078$ with a censoring rate of $79\%$, since $225$ patients were clinically diagnosed during the study while $853$ were not.
@@
To develop the MRS model, Long and Paulsen (2015) Long2015 began by conducting a variable selection procedure. They used random survival forests, a nonparametric technique that accommodates non-linear and interaction effects while avoiding specific distributional assumptions about $X_{i}$. Candidate models with varying sets of covariates were evaluated based on an inverse probability of censoring (IPCW)-based Brier score Park2021.
@@
$\lambda_{0}(t)\exp($ $\beta_{1}\text{DCL}1_{i}+\beta_{2}\text{DCL}2_{i}+\beta_{3}\text{DCL}3_{i}+\beta_{4}\text{TMS}_{i}+\beta_{5}\text{Color}_{i}+\beta_{6}\text{Word}_{i}+\beta_{7}\text{Interference}_{i}$
$+$ $\beta_{8}\text{SDMT}_{i}+\beta_{9}\text{CAG}_{i}+\beta_{10}\text{Age}_{i}+\beta_{11}\text{TMS}_{i}^{2}+\beta_{12}\text{CAG}_{i}\times\text{TMS}_{i}+\beta_{13}\text{CAG}_{i}\times\text{Age}_{i})$
@@
Like Langbehn and CAP, the MRS model assumes that $X$ is conditionally independent of $C$ given $\boldsymbol{Z}$. It also uniquely assumes that the hazard ratio between two patients with different covariates remains constant over time (i.e., proportional hazards). Similar to the CAP model, the linear predictor of the MRS model, ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}$, is interpretable as a log-risk score, which enables risk stratification. Moreover, the model can be estimated using the coxph() function from the survival package in R Therneau2022.
@@ 2.2.4 Model 4: Prognostic Index Normed (PIN)
Most recently, Long et al. (2017) Long2017 introduced the PIN model, which is another Cox model of time to HD diagnosis. In contrast to the Langbehn, CAP, and MRS models, the PIN model was developed using integrated data from three studies: PREDICT-HD, COHORT, and TRACK-HD. Each dataset was restricted to include only patients who met the following criteria at enrollment: (i) $\text{CAG}\geq 36$ and (ii) $\text{DCL}<4$. These criteria, which are identical to those for CAP and MRS, left a total sample size of $1421$ patients, with a right-censoring rate of $77\%$.
Development of the PIN model began with a variable selection procedure. Long et al. (2017) Long2017 fit ten pre-planned Cox regression models (drawing from their earlier HD research) with time to diagnosis $X_{i}$ as the outcome and different combinations of seven covariates: DCL, TMS, Stroop, SDMT, CAG, Age, and $\textrm{CAP}=\textrm{Age}(\textrm{CAG}-34)$. These candidate models were fit using only the PREDICT-HD data, and Harrell’s C statistic was used to evaluate each model’s performance. To choose from among the ten covariate combinations, leave-one-site-out cross validation was used. That is, the candidate models were estimated using data from all clinical sites in PREDICT-HD, omitting one, and then its predictive performance was measured via Harrell’s C using data from the held-out site.
@@
Since PIN is also a Cox model, it shares several properties with MRS. First, it makes similar assumptions: (i) $X$ is conditionally independent of $C$ given $\boldsymbol{Z}$ and (ii) proportional hazards. Similar to the CAP and MRS models, ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}$ from the PIN model is interpretable as a log-risk score and can be used to stratify patients by risk. The model can be estimated using the coxph() function from the survival package in R Therneau2022.
@@ 2.3 Modeling Trade-Offs
Table 3: Summary of existing models for time to diagnosis of Huntington disease: Langbehn, CAG-Age Product (CAP), Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN).
@@ 2.4 Metrics for Evaluating Risk Stratification
2.4.1 Metric 1: Uno’s Concordance (C) Statistic
@@
To measure predictive model performance with heavily censored data, Uno et al. (2011) Uno2011 introduced a novel concordance index that incorporates IPCW to make use of the full dataset. Specifically, Uno’s C statistic weights each pair ($i$, $j$) of uncensored observations ($i=1,\dots,n$; $j=1,\dots,n$; $i\neq j$) based on the estimated probability that the patient with the smaller observed time $W$ (time to diagnosis or censoring) remains uncensored up to $\min(W_{i},W_{j})$. These weights are calculated from the Kaplan-Meier survival estimator for $C$, denoted by $\widehat{G}(c)=\widehat{\Pr}(C>c)$.
@@
The central quantity in Uno’s C statistic is ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$, which denotes patient $i$’s log-risk score given covariates $\boldsymbol{Z}_{i}$ and parameters $\boldsymbol{\beta}$}. The Langbehn model does not produce a log-risk score, but we can use the conditional probability from (2) as a proxy when calculating its Uno’s C statistic. Letting $\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{i}$ denote the estimated log-risk score for patient $i$ given parameter estimates $\hat{{\boldsymbol{\beta}}}$ from a particular model, Uno’s C statistic is calculated as follows:
@@
$C_{\text{U}}(\hat{{\boldsymbol{\beta}}},\tau)=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\Delta_{i}\left\{\widehat{G}(W_{i})^{-2}\right\}\text{I}\left(W_{i}<W_{j},W_{i}<\tau\right){\rm I}\left(\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{i}>\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{j}\right)}{\sum_{i=1}^{n}\sum_{j\neq i}\Delta_{i}\left\{\widehat{G}(W_{i})^{-2}\right\}\text{I}\left(W_{i}<W_{j},W_{i}<\tau\right)},$ (3)
@@
where $\tau$ is the time horizon of interest such that $\widehat{G}(\tau)>0$. We calculate $C_{\text{U}}(\hat{{\boldsymbol{\beta}}},\tau)$ for a range of $\tau$ values to compare the models across several time horizons. Available software includes the UnoC() function from the survAUC package in R Potapov2024.
@@
Formally, Uno’s C statistic estimates the probability that a randomly selected patient who was diagnosed before time $\tau$ and earlier than another patient is assigned a higher log-risk score than that other patient. Uno’s C estimates close to $1$ indicate strong risk stratification, whereas values below $0.5$ indicate model performance that is worse than a coin flip. Uno’s C statistic makes greater use of the available data by incorporating $\widehat{G}(c)$ as weights, making it more robust to heavy censoring than Harrell’s C. Although (3) still discards some “incompatible pairs” (i.e., any $[i,j]$ with both patients censored, or $W_{i}<W_{j}$ but $\Delta_{i}=0$), these pairs still indirectly contribute to Uno’s C statistic by informing $\widehat{G}(c)$.
@@ 2.4.2 Metric 2: Receiver Operating Characteristic (ROC) Curves
In survival analysis, time-dependent receiver operating characteristic (ROC) curves are also used to evaluate a model’s risk stratification ability at specific time points. These curves plot a model’s sensitivity (true positive rate) on the y-axis against $(1-\text{specificity})$ (false positive rate) on the x-axis at a given time $t$, indicating how well it distinguishes between patients who are diagnosed by time $t$ and those who are not. The true positive rate ($\text{TPR}(q,t)$) and false positive rate ($\text{FPR}(q,t)$) for threshold $q$ at time $t$ are defined as $\text{TPR}(q,t)=\Pr(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\mid X_{i}\leq t)$ and $\text{FPR}(q,t)=\Pr(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\mid X_{i}>t)$, where $\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ continues to denote the estimated log-risk score for patient $i$ from a given model. These quantities represent the proportions of patients who are correctly or incorrectly classified as high risk (i.e., $\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q$) at time $t$, based on whether they have been diagnosed by that time (i.e., $X_{i}\leq t$).
In practice, we calculate $\text{TPR}(q,t)$ and $\text{FPR}(q,t)$ over the range of thresholds $q$ from $\min(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{1},\ldots,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{n})$ to $\max(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{1},\ldots,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{n})$. We obtain the ROC curve for fixed time $t$ by plotting the pairs $\{\text{FPR}(q,t),\text{TPR}(q,t)\}$ over the range of $q$. By evaluating ROC curves at multiple time points (e.g., at 3, 5, or 10 years), we can assess a model’s performance across the follow-up period.
In the absence of censoring, we could estimate $\text{TPR}(q,t)$ and $\text{FPR}(q,t)$ using sample proportions as $\widetilde{\text{TPR}}(q,t)=\sum_{i=1}^{n}\text{I}(X_{i}\leq t,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)/\sum_{i=1}^{n}\text{I}(X_{i}\leq t)$ and $\widetilde{\text{FPR}}(q,t)=\sum_{i=1}^{n}\text{I}(X_{i}>t,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)/\sum_{i=1}^{n}\text{I}(X_{i}>t)$. However, when $X_{i}$ is heavily right-censored, as in observational data of HD, these estimates can be biased. To address this bias, Heagerty et al. (2000) heagerty2000time proposed censoring-adjusted estimators:
@@
$\widehat{\text{FPR}}(q,t)$ $=1-\frac{\widehat{S}\left(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q\right)\cdot\widehat{\Pr}\left(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q\right)}{\widehat{S}(t)},$
where $\widehat{S}(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)$ and $\widehat{S}(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q)$ denote the conditional Kaplan-Meier survival estimators for $X$ among patients whose log-risk scores exceed or fall below $q$, respectively; $\widehat{\Pr}({\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)$ and $\widehat{\Pr}({\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q)$ denote the sample proportions in these two groups; and $\widehat{S}(t)$ the marginal Kaplan-Meier survival estimator for $X$ evaluated at $t$.
To summarize a model’s risk stratification ability based on ROC analysis, we can calculate the area under the ROC curve (AUC). AUC estimates take on values in the range $[0,1]$, with larger values indicating stronger risk stratification and values <0.5<0.5 indicating performance worse than a coin flip. To compare the four models, we use the AUC of the Kaplan-Meier-adjusted ROC curves, specifically.

### v2

@@ 2.2.2 Model 2: CAG-Age Product (CAP)
The CAP model was introduced by Zhang et al. (2011) Zhang et al. (2011) to predict a patient’s time to diagnosis given their CAG repeats and current age. It was developed using data from the PREDICT-HD study, restricted to patients with (i) a CAG repeats $\geq 36$ and (ii) DCL <4<4 at study entry. That is, patients needed to be at-risk for Huntington disease but not yet diagnosed at enrollment. After applying these criteria, there were $730$ patients. During the study, $137$ of these patients were clinically diagnosed, while the remaining $593$ ($81\%$) were not and thus had right-censored times of diagnosis.
@@
The CAP model is defined $\hbox{log}(X_{i})=\beta_{0}+\beta_{1}\text{Age}_{i}(\text{CAG}_{i}+\beta_{2})+\epsilon,$ where ${\boldsymbol{\beta}}=(\beta_{0},\beta_{1},\beta_{2})$ denote the intercept and effects of $\text{Age}_{i}\times\text{CAG}_{i}$ and $\text{Age}_{i}$. The error term $\epsilon$ is assumed to follow a logistic distribution with location =0=0 and scale $\sigma$. The linear predictor, $\beta_{0}+\beta_{1}\text{Age}_{i}(\text{CAG}_{i}+\beta_{2})$, can be interpreted as a log-risk score. The parameters ($\boldsymbol{\beta}$}, $\sigma$) are estimated via MLE, with a likelihood function that accounts for right censoring. Time to diagnosis $X_{i}$ is assumed to be conditionally independent of $C_{i}$ given CAG repeats and current age. The CAP model can be estimated using the survreg() function from the survival package in R with option dist = "loglogistic" (Therneau, 2022).
@@ 2.2.3 Model 3: Multivariate Risk Score (MRS)
The MRS model was developed with similar methods and intended applications to the CAP model. Designed with observational studies in mind, the MRS model predicts the time from enrollment to HD diagnosis (Long and Paulsen, 2015). However, compared to Langbehn and CAP, MRS was developed using a much broader set of motor, imaging, cognitive, functional, psychiatric, and demographic covariates.
The MRS model was developed using data from PREDICT-HD, subsetting to patients with (i) CAG repeats $\geq 36$ and (ii) DCL <4<4 at baseline. With these criteria (the same as those for the CAP model), the analytic dataset had a sample size of $1078$ with a censoring rate of $79\%$, since $225$ patients were clinically diagnosed during the study while $853$ were not.
@@
To develop the MRS model, Long et al. (2015) Long and Paulsen (2015) began by conducting a variable selection procedure. They used random survival forests, a nonparametric technique that accommodates non-linear and interaction effects while avoiding specific distributional assumptions about $X_{i}$. Candidate models with varying sets of covariates were evaluated based on an inverse probability of censoring (IPCW)-based Brier score (Park et al., 2021).
@@
$\lambda_{0}(t)\exp($ $\beta_{1}\text{DCL}1_{i}+\beta_{2}\text{DCL}2_{i}+\beta_{3}\text{DCL}3_{i}+\beta_{4}\text{TMS}_{i}+\beta_{5}\text{Color}_{i}+$
$\beta_{6}\text{Word}_{i}+\beta_{7}\text{Interference}_{i}+\beta_{8}\text{SDMT}_{i}+\beta_{9}\text{CAG}_{i}+$
$\beta_{10}\text{Age}_{i}+\beta_{11}\text{TMS}_{i}^{2}+\beta_{12}\text{CAG}_{i}\times\text{TMS}_{i}+\beta_{13}\text{CAG}_{i}\times\text{Age}_{i})$
@@
Like Langbehn and CAP, the MRS model assumes that $X$ is conditionally independent of $C$ given $\boldsymbol{Z}$. It also uniquely assumes that the hazard ratio between two patients with different covariates remains constant over time (i.e., proportional hazards). Similar to the CAP model, the linear predictor of the MRS model, ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}$, is interpretable as a log-risk score, which enables risk stratification. Moreover, the model can be estimated using the coxph() function from the survival package in R (Therneau, 2022).
@@ 2.2.4 Model 4: Prognostic Index Normed (PIN)
Most recently, Long et al. (2017) Long et al. (2017) introduced the PIN model, which is another Cox model of time to HD diagnosis. In contrast to the Langbehn, CAP, and MRS models, the PIN model was developed using integrated data from three studies: PREDICT-HD, COHORT, and TRACK-HD. Each dataset was restricted to include only patients who met the following criteria at enrollment: (i) $\text{CAG}\geq 36$ and (ii) $\text{DCL}<4$. These criteria, which are identical to those for CAP and MRS, left a total sample size of $1421$ patients, with a right-censoring rate of $77\%$.
Development of the PIN model began with a variable selection procedure. Long et al. (2017) Long et al. (2017) fit ten pre-planned Cox regression models (drawing from their earlier HD research) with time to diagnosis $X_{i}$ as the outcome and different combinations of seven covariates: DCL, TMS, Stroop, SDMT, CAG, Age, and $\textrm{CAP}=\textrm{Age}(\textrm{CAG}-34)$. These candidate models were fit using only the PREDICT-HD data, and Harrell’s C statistic was used to evaluate each model’s performance. To choose from among the ten covariate combinations, leave-one-site-out cross validation was used. That is, the candidate models were estimated using data from all clinical sites in PREDICT-HD, omitting one, and then its predictive performance was measured via Harrell’s C using data from the held-out site.
@@
Since PIN is also a Cox model, it shares several properties with other common models. First, it makes similar assumptions: (i) $X$ is conditionally independent of $C$ given $\boldsymbol{Z}$ and (ii) proportional hazards. Second, ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}$ from the PIN model is interpretable as a log-risk score and can be used to stratify patients by risk. The model can be estimated using the coxph() function from the survival package in R (Therneau, 2022).
@@ 2.3 Modeling Trade-Offs
Table 2: Summary of existing models for time to diagnosis of Huntington disease: Langbehn, CAG-Age Product (CAP), Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN). See Figures 1 and 2 for time-specific evaluation of model performance in the ENROLL-HD study.
@@
2.4 External Validation Versus Model Updating
In practice, it would be simplest to apply a published model “as-is” to new data, and treat the predictions as established equations. For example, we could predict with $\hbox{log}(\widehat{X}_{i})=\hat{\beta}_{0}+\hat{\beta}_{1}\textrm{Age}_{i}(\textrm{CAG}_{i}+\hat{\beta}_{2})$ following the CAP model where the estimates $\hat{\beta}_{0},\hat{\beta}_{1},\hat{\beta}_{2}$ were estimated using the PREDICT-HD study and can be found in Zhange et al. (2011) Zhang et al. (2011) However, since the models thus far have only been evaluated on the same data used to train them (i.e., subject to double-dipping), it is unclear how well they would perform in a new cohort of HD patients. External validation, wherein the published coefficient estimates (based on the original training data) are applied to new testing data (ENROLL-HD), is needed to evaluate the transportability of the four popular HD models (Moons et al., 2012).
Moreover, the models were all trained on HD cohorts that ended more than $10$ years ago (Table 1), whereas many more recent studies, like ENROLL-HD, are still ongoing. Therefore, we also considered model updating, wherein the predictors and distributions (where applicable) of all models were unchanged but the corresponding coefficients were re-estimated (updated) using the ENROLL-HD data. We used $5$-fold cross-validation to evaluate the performance of the updated models, pooling the metrics from Section 2.5 across folds by taking the means. These pooled metrics should be more representative of performance in other HD studies than those obtained from a single fit, since the latter would introduce double-dipping by training and testing on ENROLL-HD.
@@ 2.5 Metrics for Evaluating Risk Stratification
We considered two censoring-adjusted metrics to evaluate the models’ risk stratification ability. First, Uno’s Concordance (Uno’s C) statistic assesses how well the models rank pairs of patients in terms of log-risk scores. Second, Kaplan-Meier-adjusted receiver operating characteristic (ROC) curves and their corresponding areas under the curve (AUC) assess how well the models separate the group of patients diagnosed by a particular time $\tau$ from the group undiagnosed by time $\tau$.
2.5.1 Metric 1: Uno’s Concordance (Uno’s C) Statistic
@@
To measure predictive model performance with heavily censored data, Uno et al. (2011) Uno et al. (2011) introduced a novel concordance index that incorporates IPCW to make use of the full dataset. Specifically, Uno’s C statistic weights each pair ($i$, $j$) of uncensored observations ($i=1,\dots,n$; $j=1,\dots,n$; $i\neq j$) based on the estimated probability that the patient with the smaller observed time $W$ (time to diagnosis or censoring) remains uncensored up to $\min(W_{i},W_{j})$. These weights are calculated from the Kaplan-Meier survival estimator for $C$, denoted by $\widehat{G}(c)=\widehat{\Pr}(C>c)$.
@@
The central quantity in Uno’s C statistic is ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$, which denotes patient $i$’s log-risk score given covariates $\boldsymbol{Z}_{i}$ and parameters $\boldsymbol{\beta}$}. The Langbehn model does not produce a log-risk score, but we can use the conditional probability from (2) as a proxy when calculating its Uno’s C statistic. Letting $\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{i}$ denote the estimated log-risk score for patient $i$ given parameter estimates $\hat{{\boldsymbol{\beta}}}$ from a particular model, Uno’s C statistic is calculated as
@@
$C_{\text{U}}(\hat{{\boldsymbol{\beta}}},\tau)=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\Delta_{i}\left\{\widehat{G}(W_{i})^{-2}\right\}\text{I}\left(W_{i}<W_{j},W_{i}<\tau\right){\rm I}\left(\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{i}>\hat{{\boldsymbol{\beta}}}^{\rm T}\boldsymbol{Z}_{j}\right)}{\sum_{i=1}^{n}\sum_{j\neq i}\Delta_{i}\left\{\widehat{G}(W_{i})^{-2}\right\}\text{I}\left(W_{i}<W_{j},W_{i}<\tau\right)},$
@@
where $\tau$ is the time horizon of interest such that $\widehat{G}(\tau)>0$. We calculate $C_{\text{U}}(\hat{{\boldsymbol{\beta}}},\tau)$ for a range of $\tau$ values to compare the models across several time horizons. Available software includes the UnoC() function from the survAUC package in R (Potapov et al., 2024).
@@
Formally, Uno’s C statistic estimates the probability that a randomly selected patient who was diagnosed before time $\tau$ and earlier than another patient is assigned a higher log-risk score than that other patient. Uno’s C estimates close to $1$ indicate strong risk stratification, whereas values below $0.5$ indicate model performance that is worse than a coin flip. Uno’s C statistic makes greater use of the available data by incorporating $\widehat{G}(c)$ as weights, making it more robust to heavy censoring than Harrell’s C. Although Uno’s C still discards some “incompatible pairs” (i.e., any $[i,j]$ with both patients censored, or $W_{i}<W_{j}$ but $\Delta_{i}=0$), these pairs still indirectly contribute to Uno’s C statistic by informing $\widehat{G}(c)$.
@@ 2.5.2 Metric 2: Receiver Operating Characteristic (ROC) Curves
In survival analysis, time-dependent ROC curves are also used to evaluate a model’s risk stratification ability at specific time points. These curves plot a model’s sensitivity (true positive rate) on the y-axis against $(1-\text{specificity})$ (false positive rate) on the x-axis at a given time $t$, indicating how well it distinguishes between patients who are diagnosed by time $t$ and those who are not. The true positive rate ($\text{TPR}(q,t)$) and false positive rate ($\text{FPR}(q,t)$) for threshold $q$ at time $t$ are defined as $\text{TPR}(q,t)=\Pr(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\mid X_{i}\leq t)$ and $\text{FPR}(q,t)=\Pr(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\mid X_{i}>t)$, where $\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ continues to denote the estimated log-risk score for patient $i$ from a given model. These quantities represent the proportions of patients who are correctly or incorrectly classified as high risk (i.e., $\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q$) at time $t$, based on whether they have been diagnosed by that time (i.e., $X_{i}\leq t$).
In practice, we calculate $\text{TPR}(q,t)$ and $\text{FPR}(q,t)$ at thresholds $q$ ranging from $\min(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{1},\ldots,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{n})$ to $\max(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{1},\ldots,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{n})$. We obtain the ROC curve at time $t$ by plotting the $\{\text{FPR}(q,t),\text{TPR}(q,t)\}$ pairs over the range of thresholds $q$. By evaluating ROC curves at multiple time points (e.g., at $3$, 55, or $10$ years), we can assess a model’s performance across the follow-up period.
In the absence of censoring, we could estimate $\text{TPR}(q,t)$ and $\text{FPR}(q,t)$ using sample proportions as
$\widetilde{\text{TPR}}(q,t)$ $=\sum_{i=1}^{n}\text{I}\left(X_{i}\leq t,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\right)/\sum_{i=1}^{n}\text{I}(X_{i}\leq t),\;\text{and}$
$\widetilde{\text{FPR}}(q,t)$ $=\sum_{i=1}^{n}\text{I}\left(X_{i}>t,\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q\right)/\sum_{i=1}^{n}\text{I}(X_{i}>t).$
When censoring is present, the same formulas could be applied to only the subset of uncensored observations, essentially leading to complete-case estimates of the TPR and FPR. However, such metrics could only describe the model’s risk stratification ability among diagnosed patients. Particularly in observational data of HD, when $X_{i}$ is heavily right-censored, $\widetilde{\text{TPR}}(q,t)$ and $\widetilde{\text{FPR}}(q,t)$ would not comprehensively describe model performance across the entire sample.
Heagerty et al. (2000) Heagerty et al. (2000) proposed censoring-adjusted estimators that incorporate censored observations and describe model performance for all patients:
@@
$\widehat{\text{FPR}}(q,t)$ $=1-\frac{\widehat{S}\left(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q\right)\widehat{\Pr}\left(\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q\right)}{\widehat{S}(t)},$
where $\widehat{S}(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)$ and $\widehat{S}(t\mid\widehat{\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q)$ denote the conditional Kaplan-Meier survival estimators for $X$ among patients whose log-risk scores exceed or fall below $q$, respectively; $\widehat{\Pr}({\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}>q)$ and $\widehat{\Pr}({\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}\leq q)$ denote the sample proportions in these two groups; and $\widehat{S}(t)$ denotes the marginal Kaplan-Meier survival estimator for $X$ evaluated at $t$.
To summarize a model’s risk stratification ability based on ROC analysis, we can calculate the AUC. AUC estimates take on values in the range $[0,1]$, with larger values indicating stronger risk stratification and values <0.5<0.5 indicating performance worse than a coin flip. To compare the four models, we use the AUC of the Kaplan-Meier-adjusted ROC curves, specifically.